OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x^6/x^2. Explain in your own words how you can simplify x^6/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

part 1 not clear..

OpenStudy (anonymous):

simplify it?

OpenStudy (anonymous):

@Krishnadas sorry i dont understand it myself i really need help

OpenStudy (anonymous):

I think it means to simplify that fraction

OpenStudy (anonymous):

x^6=X^3.x^2 x^6/x^2=x^3

OpenStudy (anonymous):

@Krishnadas what about part 2

OpenStudy (anonymous):

idk what they meant..:(

OpenStudy (anonymous):

thank you anyway @Krishnadas

OpenStudy (anonymous):

okay..

jimthompson5910 (jim_thompson5910):

x^6/x^2 is actually x^4 since 6-2 = 4